Match the IR spectrum to one of the following thre
Answer #1

2-methy -1-octene     An ALKENE

Butyl methyl ether   An ETHER

1-Pentanol               An ALCOHOL

Aekene C-H No 3o20 -308o 675 Iovo an alkens e 360-3640N
Answer #2
Concepts and reason

Infrared spectroscopy is used to determine the functional groups present in a compound. Infrared bands observed when there is change in dipole moment occurs between the atoms. Infrared bands describe about the bond stretches causes due to the dipole moment present in the molecule.

Fundamentals

Infrared spectroscopy:

Infrared bands appear at the certain absorption frequencies.

wavenumber(or)absorbanceisνν=12πckμwhereμ=reducedmassc=speedoflightk=springconstntofthebond\begin{array}{l}\\{\rm{wave}}\,{\rm{number}}\,\left( {{\rm{or}}} \right)\,{\rm{absorbance}}\,\,{\rm{is}}\,\,{\rm{\nu }}\\\\{\rm{\nu }}\,{\rm{ = }}\,\frac{{\rm{1}}}{{{\rm{2\pi c}}}}\sqrt {\frac{{\rm{k}}}{{\rm{\mu }}}} \\\\{\rm{where}}\\\\{\rm{\mu = }}\,{\rm{reduced}}\,{\rm{mass}}\\\\{\rm{c }}\,{\rm{ = }}\,{\rm{speed}}\,{\rm{of}}\,{\rm{light}}\\\\{\rm{k = }}\,{\rm{spring}}\,{\rm{constnt}}\,{\rm{of}}\,{\rm{the}}\,{\rm{bond}}\\\end{array}

Reduced mass (μ)\left( {\rm{\mu }} \right) and spring constant (k)\left( {\rm{k}} \right) of a bond changes with respect to the functional groups. Stronger bonds with lighter atoms are having higher frequency range.

Given compound

2-methyl-1-octene

This is unsaturated hydrocarbon belongs to alkene (C=C)\left( {{\rm{C = C}}} \right) type.

Functional group present in the compound is “ene”

IR range: (16001500)cm1\left( {{\rm{1600 - 1500}}} \right){\rm{c}}{{\rm{m}}^{{\rm{ - 1}}}}

1-pentanol

This compound belongs to alcohol category.

Functional group present in the molecule is alcohol (OH).\left( { - {\rm{OH}}} \right).

IR range: (36503590)cm1\left( {{\rm{3650 - 3590}}} \right){\rm{c}}{{\rm{m}}^{{\rm{ - 1}}}}

Butyl methyl ether

This compound belongs to ether category.

Functional group present in the molecule is ether (COC).\left( {{\rm{C}} - {\rm{O}} - {\rm{C}}} \right).

IR range: (12001080)cm1\left( {{\rm{1200 - 1080}}} \right){\rm{c}}{{\rm{m}}^{{\rm{ - 1}}}}

Given IR spectrum

2.6 2.8 3
3.5
4 4.5
wavelength, micrometers
5 5.5 6 7 8
9 10
11
12 13 14 1516
percent transmittance
800
600
3800 3400 3000 26

From this IR spectrum, it is clear only Butyl methyl ether band merges with the spectrum. In the spectrum, there is a sharp peak between (11001200)cm1\left( {{\rm{1100 - 1200}}} \right){\rm{c}}{{\rm{m}}^{{\rm{ - 1}}}} .

In the spectrum, there are no other sharp peaks in the range of alcohol functional group and for carbon –carbon double bond.

Therefore, this spectrum is for Butyl methyl ether compound.

Ans:

02-methyl-1-octene
0
1-pentanol
butyl methyl ether